### v1

# Faithful Group Shapley Value

Abstract
Data Shapley is an important tool for data valuation, which quantifies the contribution of individual data points to machine learning models. In practice, group-level data valuation is desirable when data providers contribute data in batch. However, we identify that existing group-level extensions of Data Shapley are vulnerable to shell company attacks, where strategic group splitting can unfairly inflate valuations. We propose Faithful Group Shapley Value (FGSV) that uniquely defends against such attacks. Building on original mathematical insights, we develop a provably fast and accurate approximation algorithm for computing FGSV. Empirical experiments demonstrate that our algorithm significantly outperforms state-of-the-art methods in computational efficiency and approximation accuracy, while ensuring faithful group-level valuation.
1 Introduction
As data become increasingly crucial in modern machine learning, quantifying its value has significant implications for fair compensation and data market design [29, 39]. The Shapley value, a foundational concept from cooperative game theory [30], stands out as the unique valuation method satisfying four desirable axioms for fair data valuation [6, 13]. Consequently, it has been applied to diverse applications, including collaborative intelligence [22, 25] and copyright compensation [35], and quantifying feature importance in explainable AI [26].
[…]
Many real-world scenarios demand evaluation of data sets instead of individual data points. In applications like data marketplaces [39] or compensation allocation for generative AI [4, 35], data are contributed by owners who possess entire datasets, making group-level data valuation a natural choice. A more technical motivation is that individual-level Shapley values are often computationally challenging to approximate for big data; while group-level valuation is much faster and can provide useful insights [16]. In explainable AI, group-level feature importance reportedly provides more robust and interpretable valuation [32, 3].
Prior works extend the Shapley framework to group data valuation by treating each pre-defined group as an atomic unit and then applying the standard Shapley value formulation [14, 20, 35, 32], which we refer to as Group Shapley Value (GSV). However, we demonstrate—both theoretically and empirically—that this approach suffers from susceptibility to strategic manipulation through partitioning. Consider the valuation of a fixed data group $A$. As shown in Figure 1, partitioning the remaining data points into smaller subgroups reduces the GSV for $A$, despite no change in data content. In other words, unfair players may exploit the loophole in GSV by splitting their data among puppy subsidiaries, a vulnerability we coin as shell company attack.
To address this issue, we propose a faithfulness axiom for group data valuation, such that the total valuation of the same set of data remains unchanged, regardless of how others are subdivided. Based on this, we introduce the Faithful Group Shapley Value (FGSV) for a group as the sum of individual Data Shapley values of its members, and prove that it uniquely satisfies the set of axioms that are desirable for faithful group data valuation. Figure 1 demonstrates that FGSV effectively defends against the shell company attack.
While FGSV safeguards fairness, its exact computation requires combinatorial computation of individual Shapley values. Although numerous approximation algorithms for individual Shapley values have been developed [2, 6, 12, 34, 26, 3, 38, 24, 27], simply applying them and summing up results in each group is slow for large groups and may compound approximation errors. Importantly, we proved that a small subset of terms dominate in FGSV’s formula. As a result, we developed an original algorithm that directly approximates FGSV fast and accurately. Our numerical experiments demonstrate its computational efficiency and approximation accuracy against summing up individual Shapley values computed by state-of-the-art (SOTA) methods. We also applied FGSV to fair copyright compensation on Stable Diffusion models for image generation.
2 Preliminaries: individual and group Shapley values
Individual Shapley value (SV). Let $\mathcal{D}=\{z_{1},\ldots,z_{n}\}\in\mathcal{Z}^{n}$ denote the training data, where $n:=|\mathcal{D}|$. Let $S\subseteq[n]$ be an index set of cardinality $s:=|S|$, and write ${\cal S}:=\{z_{i}:i\in S\}$. A utility function $U({\cal S})$ assigns a performance score (such as classification accuracy) to method trained based on data ${\cal S}$. To simplify notation, we may refer to “$U({\cal S})$” by index as “$U(S)$” and write them interchangeably. The individual Shapley value of data point $i$ is defined as
$\operatorname{SV}(i):=\sum_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}\{U(S\cup\{i\})-U(S)\}.$ (1)
The widespread adoption of Shapley value stems from its strong theoretical foundation that it uniquely satisfies four desired axioms, called null player, symmetry, linearity, and efficiency. We will review these principles as part of our axiomatization for group Shapley value in Definition 2.
Group Shapley value (GSV). As mentioned in Section 2, most group-level Shapley values [14, 35, 32] adopt the so-called group-as-individual (GaI) approach. Suppose the entire data set is partitioned into $(K+1)$ disjoint groups: $\mathcal{D}=S_{0}\cup S_{1}\cup\cdots\cup S_{K}$, in which, $S_{0}\subseteq[n]$ is the group we aim to evaluate. The group Shapley value (GSV) of $S_{0}$, as in [14, 35, 32], is defined as
$\operatorname{GSV}(S_{0})=$ $~{}\sum_{I\subseteq\{1,\dots,K\}}\frac{|I|!(K-|I|)!}{(K+1)!}\Big{\{}U\Big{(}\big{\{}\cup_{k\in I}S_{k}\big{\}}\cup S_{0}\Big{)}-U\Big{(}\cup_{k\in I}S_{k}\Big{)}\Big{\}}.$ (2)
Clearly, (2) is a “set version” of (1). Thus GSV also satisfies the four axioms for SV at the group level. However, it is questionable whether these axioms, originally formulated for individual-level valuation, are appropriate for groups.
Fairness issue with GSV. The GSV valuation of $S_{0}$ as in (2) can be impacted by how the rest of the data are grouped. A serious consequence is that GSV is prone to the shell company attack: splitting a group into smaller subgroups will increase the total valuation earned by the same set of data points. We have seen an illustration in Figure 1 in Section 1. Now we theoretically formalize this observation.
Proposition 1 (Shell company attack).
Suppose in $\mathcal{D}=\{z_{1},\dots,z_{n}\}$ are an i.i.d. sample. Let $\bar{U}(s)=\mathbb{E}_{S\sim\mathcal{P}^{s}}[U(S)]$ denote the expected utility for data $S$ (which only depends on $s=|S|$). Now we split a group $S_{k}$ into two non-empty subgroups $S_{k}^{\prime}$ and $S_{k}^{\prime\prime}$ (i.e., $S_{k}^{\prime}\cup S_{k}^{\prime\prime}=S_{k}$, $S_{k}^{\prime}\cap S_{k}^{\prime\prime}=\emptyset$, and $S_{k}^{\prime},S_{k}^{\prime\prime}\neq\emptyset$), If $\bar{U}(s)$ satisfies a prudence condition:
$\Delta_{s}^{3}\bar{U}(s):=\bar{U}(s+3)-3\bar{U}(s+2)+3\bar{U}(s+1)-\bar{U}(s)>0,$ (3)
then
$\mathbb{E}[\operatorname{GSV}(S_{k})]<\mathbb{E}[\operatorname{GSV}(S_{k}^{\prime})]+\mathbb{E}[\operatorname{GSV}(S_{k}^{\prime\prime})].$ (4)
Condition (3) is a familiar concept in economics, characterizing risk-aversion behaviors under uncertainty [7]. It is also observed in machine learning, where the performance of a learning method saturates as the same type of data repeatedly come in [17, 10].
3 Our method
3.1 Faithfulness axiom and Faithful Group Shapley Value (FGSV)
Motivated by the faithfulness issue, we introduce a set of axioms that formalize desirable principles for faithful group data valuation. We first axiomatize group-level data valuation that generalizes the Shapley’s four axioms for individual valuation.
Definition 1 (Group data valuation).
For data $\mathcal{D}=\{z_{1},\dots,z_{n}\}$, utility $U$ and partition $\Pi=\{S_{1},\dots,S_{K}\}$ such that $S_{k_{1}}\cap S_{k_{2}}=\emptyset,\forall 1\leq k_{1}<k_{2}\leq n$ and $S_{1}\cup\cdots\cup S_{K}=[n]$, a group data valuation method $\nu_{U,\mathcal{D},\Pi}(\cdot):\Pi\to\mathbb{R}$ is a mapping that assigns a real-valued score to each $S_{k}\in\Pi$.
Next, we propose a set of axioms that a faithful group data valuation method should satisfy.
Definition 2 (Axioms for faithful group data valuation).
A group data valuation $\nu_{U,\mathcal{D},\Pi}$ is called faithful if it satisfies the following axioms for any dataset $\mathcal{D}$, utility $U$, and partition $\Pi$:
  1. 1.
    Null player: For any $S\in\Pi$, if every subset $S^{\prime}\subseteq S$ satisfies $U(S^{\prime\prime}\cup S^{\prime})=U(S^{\prime\prime})$ for all $S^{\prime\prime}\subseteq[n]\setminus S^{\prime}$, then $\nu_{U,\mathcal{D},\Pi}(S)=0$.
  2. 2.
    Symmetry: For any $S_{1},S_{2}\in\Pi$, $|S_{1}|=|S_{2}|$, if there is a bijection $\sigma:S_{1}\to S_{2}$, s.t. $U(S^{\prime\prime}\cup S^{\prime})=U(S^{\prime\prime}\cup\sigma(S^{\prime}))$ for all $S^{\prime}\subseteq S_{1}$ and $S^{\prime\prime}\subseteq[n]\setminus(S^{\prime}\cup\sigma(S^{\prime}))$, then $\nu_{U,\mathcal{D},\Pi}(S_{1})=\nu_{U,\mathcal{D},\Pi}(S_{2})$.
  3. 3.
    Linearity: For any utility functions $U_{1},U_{2}$ and scalars $\alpha_{1},\alpha_{2}\in\mathbb{R}$, we have for all $S\in\Pi$ $\nu_{\alpha_{1}U_{1}+\alpha_{2}U_{2},\mathcal{D},\Pi}(S)=\alpha_{1}\nu_{U_{1},\mathcal{D},\Pi}(S)+\alpha_{2}\nu_{U_{2},\mathcal{D},\Pi}(S)$.
  4. 4.
    Efficiency: For any partition $\Pi=\{S_{1},\dots,S_{K}\}$, we have $\sum_{k=1}^{K}\nu_{U,\mathcal{D},\Pi}(S_{k})=U([n])$.
  5. 5.
    Faithfulness: For any group $S\in\Pi_{1}\cap\Pi_{2}$, we have $\nu_{U,\mathcal{D},\Pi_{1}}(S)=\nu_{U,\mathcal{D},\Pi_{2}}(S)$.
Notably, our null player and symmetry require group valuation to faithfully reflect the contributions of its individual members. They are strictly weaker assumptions than their GaI counterparts. For example, a group assigned zero value in GaI can receive nonzero value under our axioms if some of its members have marginal contributions, but not the other way around. Linearity and efficiency are standard group-level extensions. Our newly introduced Axiom 5 requires that a group’s value is determined only on its own competitive merit and rules out the unfair competition means of shell company attack.
It turns out that there is a unique group data valuation method that can satisfy all axioms in Definition 2: simply add up all member’s individual Shapley values.
Theorem 1.
The only group data valuation method $\nu_{U,\mathcal{D},\Pi}$ that satisfies all axioms in Definition 2 is $\nu_{U,\mathcal{D},\Pi}(S)=\sum_{i\in S}\operatorname{SV}(i),$ where $\operatorname{SV}(i)$ is the individual Shapley value defined in (1).
In view of Theorem 1, we propose Faithful Group Shapley Value (FGSV):
$\operatorname{FGSV}(S_{0}):=\sum_{i\in S_{0}}\operatorname{SV}(i).$ (5)
3.2 Fast and accurate approximation algorithm for FGSV
The exact evaluation of FGSV requires combinatorial computation. To develop a feasible approximation method, we make a series of key mathematical observations leading to an efficient algorithm.
To start, from the definition of FGSV (5) and (1), we see that $\operatorname{FGSV}(S_{0})$ is a complicated linear combination of $U(S)$ terms, where $S$ ranges over all subsets of $[n]$. Therefore, the first step towards simplification is to discover the pattern in the coefficient in front of each $U(S)$ term.
Key observation 1.
In $\operatorname{FGSV}(S_{0})$, the coefficient of $U(S)$ depends on $S$ and $S_{0}$ only through the tuple $(s_{1},s,s_{0}),$ where recall that $s_{0}:=|S_{0}|$ and $s:=|S|$, and define $s_{1}:=|S_{0}\cap S|$.
In other words, any two terms $U(S)$ and $U(S^{\prime})$ with $|S|=|S^{\prime}|$ and $|S_{0}\cap S|=|S_{0}\cap S^{\prime}|$ share the same coefficient in $\operatorname{FGSV}(S_{0})$. This motivates us to aggregate these terms in our analysis. Let $\mathscr{A}_{s,s_{1}}:=\{S:|S|=s,|S\cap S_{0}|=s_{1}\}$ collect all $S$’es with the same $(s,s_{1})$ configuration, and define
$\mu\left(\frac{s_{1}}{s};s,s_{0},n\right):=$ $~{}\frac{\sum_{S\in{\cal A}_{s,s_{1}}}U(S)}{|{\cal A}_{s,s_{1}}|}=\frac{\sum_{S:|S|=s,|S\cap S_{0}|=s_{1}}U(S)}{\binom{s_{0}}{s_{1}}\binom{n-s_{0}}{s-s_{1}}}.$ (6)
When sampling a subset $S\subseteq[n]$ of size $s$ without replacement, the size of $S\cap S_{0}$, which we now denote as the boldfaced $\bm{s_{1}}$ to emphasize its randomness, follows a hypergeometric distribution: $\mathbb{P}(\bm{s_{1}}=s_{1})=\binom{s_{0}}{s_{1}}\binom{n-s_{0}}{s-s_{1}}/\binom{n}{s}.$ Using this fact, we can re-express FGSV in terms of $\mu$.
Lemma 1.
Let $\bm{s_{1}}\sim\mathcal{HG}(n,s_{0},s)$. We can rewrite FGSV as
$\operatorname{FGSV}(S_{0})=$ $~{}\frac{s_{0}}{n}\left[U([n])-U(\varnothing)\right]+\sum_{s=1}^{n-1}{\cal T}(s),$ (7)
where
${\cal T}(s):=$ $~{}\mathbb{E}_{\bm{s_{1}}\sim\mathcal{HG}(n,s_{0},s)}\left[\frac{n}{n-s}\left(\frac{\bm{s_{1}}}{s}-\frac{s_{0}}{n}\right)\mu\left(\frac{\bm{s_{1}}}{s};s,s_{0},n\right)\right].$ (8)
Instead of directly estimating ${\cal T}(s)$ via Monte Carlo, we discover two key observations that deepen our understandings of (7) and (8), building on which, we can greatly reduce computational cost.
Key observation 2.
The probability $\mathbb{P}(\bm{s_{1}}=s_{1})$ decays exponentially in $|s_{1}-ss_{0}/n|$.
Key observation 2 implies that ${\cal T}(s)$ is dominated by the values of $\bm{s_{1}}$ around its mean $\mathbb{E}[\bm{s_{1}}]=ss_{0}/n$. To deepen our understanding of ${\cal T}(s)$, for this moment, we informally deem $\mu$ as a smooth function of the continuous variable $s_{1}/s$, with formal characterization provided later. Then, applying a “Taylor expansion” of $\mu\left(s_{1}/s;s,s_{0},n\right)$ around $s_{1}/s=s_{0}/n$ leads to the following intuition.
Key observation 3 (Informal).
${\cal T}(s)\approx s^{-1}(s_{0}/n)(1-s_{0}/n)\mu^{\prime}(s_{0}/n;s,s_{0},n).$
Key observation 3 reveals that, under suitable conditions, the term ${\cal T}(s)$ can be efficiently estimated by evaluating the derivative $\mu^{\prime}\left(\cdot;s,s_{0},n\right)$ at a single point $s_{0}/n$.
Next, we formalize the above discoveries into rigorous mathematical results.
Assumption 1 (Boundedness).
For all $s\in\mathbb{N}$ and $\mathcal{S}\in\mathcal{Z}^{s}$, $|U(\mathcal{S})|\leq C$ for a universal constant $C$.
Assumption 2 (Second-order algorithmic stability of utility).
There exist constants $C>0$ and $\upsilon>0$ such that for all $s\in\mathbb{N}$, $\mathcal{S}\in\mathcal{Z}^{s}$ and $z_{1},z_{1}^{\prime},z_{2},z_{2}^{\prime}\in\mathcal{Z}$,
$\big{|}U(\mathcal{S}\cup\{z_{1},z_{1}^{\prime}\})-U(\mathcal{S}\cup\{z_{1},z_{2}\})-U(\mathcal{S}\cup\{z_{1}^{\prime},z_{2}^{\prime}\})+U(\mathcal{S}\cup\{z_{2},z_{2}^{\prime}\})\big{|}\leq Cs^{-(3/2+\upsilon)}.$
Assumption 1 is a common regularity condition, and Assumption 2 introduces a mild second-order stability requirement. In Section 3.5, we will show examples that Assumption 2 is satisfied by some commonly used utility functions.
Theorem 2.
Under Assumptions 1 and 2, for each $s\in\{1,\dots,n-1\}$, we have
$\mathcal{T}(s)=$ $~{}n/(n-1)\cdot\alpha_{0}(1-\alpha_{0})\Big{\{}\Delta\mu\big{(}s_{1}^{*}/s;s,s_{0},n\big{)}+O\big{(}s^{-(1+\upsilon)}\big{)}\Big{\}},$ (9)
where $s_{1}^{*}:=\lfloor ss_{0}/n\rfloor$, $\alpha_{0}:=s_{0}/n$, and $\Delta\mu(\frac{s_{1}}{s};s,s_{0},n):=\mu(\frac{s_{1}+1}{s};s,s_{0},n)-\mu(\frac{s_{1}}{s};s,s_{0},n)$.
Theorem 2 consolidates Key observation 3 and, moreover, shows that its approximation error decays rapidly as $s$ grows. For large $s$, we can use (9) to design the estimator for $\mathcal{T}(s)$. Yet, some extra care is needed. In principle, each $\mu(s_{1}/s;s,s_{0},n)$ can be estimated by subsampling $S$ from ${\cal A}_{s,s_{1}}$.
$\widehat{\mu}_{m}\left(\frac{s_{1}}{s};s,s_{0},n\right):=$ $~{}\frac{1}{m}\sum_{j=1}^{m}U(S^{(j)}),\quad\text{where }\{S^{(j)}\}_{j=1}^{m}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\operatorname{Uniform}(\mathscr{A}_{s,s_{1}}).$ (10)
However, estimating $\mu(\frac{s_{1}+1}{s};s,s_{0},n)$ and $\mu(\frac{s_{1}}{s};s,s_{0},n)$ separately can be statistically inefficient, as the noise from two independent Monte Carlo estimates could mask the signal in their difference—especially when the true gap is small for large $s$. Following the variance-reduction technique used in stochastic simulation [21], we propose to estimate $\Delta\mu$ directly using paired Monte Carlo terms:
$\widehat{\Delta\mu}_{m}\left(\frac{s_{1}}{s};s,s_{0},n\right):=$ $~{}\frac{1}{m}\sum_{j=1}^{m}\Big{\{}U\big{(}S^{(j)}\cup\{i_{1}^{(j)}\}\big{)}-U\big{(}S^{(j)}\cup\{i_{2}^{(j)}\}\big{)}\Big{\}},$ (11)
where the tuple $\{(S^{(j)},i_{1}^{(j)},i_{2}^{(j)})\}$ is i.i.d. sampled from $\big{\{}(S,i_{1},i_{2}):|S|=s-1,|S\cap S_{0}|=s_{1},i_{1}\in S_{0}\setminus S,i_{2}\in S_{0}^{c}\setminus S\big{\}}$.
When $s$ is small and the approximation becomes less accurate, we instead estimate ${\cal T}(s)$ via direct Monte Carlo using (8) and (10).
Algorithm 1 Approximate $\operatorname{FGSV}(S_{0})$
0:  Dataset $\mathcal{D}$, group $S_{0}$, threshold $\bar{s}$, subsample sizes $m_{1},m_{2}$.
1:  Initialize $n=|\mathcal{D}|$, $s_{0}=|S_{0}|$ and $\alpha_{0}=s_{0}/n$.
2:  for $s=1$ to $n-1$ do
3:     if $s<\bar{s}$ then
4:        Estimate $\widehat{\mu}_{m_{1}}(\frac{s_{1}}{s};s,s_{0},n)$ for each $s_{1}\in[\max\{0,s{+}s_{0}{-}n\},\min\{s,s_{0}\}]$ by Eq. (10).
5:        Compute $\widehat{\mathcal{T}}(s)$ by (8), replacing $\mu$ by $\widehat{\mu}_{m_{1}}$.
6:     else
7:        $s_{1}^{*}\leftarrow\lfloor s\alpha_{0}\rfloor$.
8:         Estimate $\widehat{\Delta\mu}_{m_{2}}(\frac{s_{1}^{*}}{s};s,s_{0},n)$ by Eq. (11).
9:        $\widehat{\mathcal{T}}(s)\leftarrow\frac{n}{n-1}\alpha_{0}(1-\alpha_{0})\cdot\widehat{\Delta\mu}_{m_{2}}(\frac{s_{1}^{*}}{s};s,s_{0},n)$.
10:     end if
11:  end for
12:  return  $\frac{s_{0}}{n}\left[U([n])-U(\varnothing)\right]+\sum_{s=1}^{n-1}\widehat{\mathcal{T}}(s)$.
We formally present our method as Algorithm 1. Later, our Theorem 3 will provide quantitative guidance on choosing this threshold for deciding whether $s$ is small or large.
3.3 Computational complexity
Following the convention, we measure computational complexity by the number of utility function evaluations required for a given approximation accuracy.
Definition 3 ($(\epsilon,\delta)$-approximation).
For a target vector $\theta\in\mathbb{R}^{d}$, an estimator $\widehat{\theta}$ is called an $(\epsilon,\delta)$-approximation, if $\mathbb{P}\big{(}\|\widehat{\theta}-\theta\|_{2}\geq\epsilon\big{)}\leq\delta.$
Our theoretical analysis for approximating $\mathcal{T}(s)$ relies on the stability of the utility function $U$. Specifically, we employ the concept of deletion stability to quantify the maximum change in the utility function when a single data point is removed [1, 9].
Definition 4 (Deletion Stability).
A utility function $U$ is $\beta(s)$-deletion stable for a non-increasing function $\beta:\mathbb{N}\to\mathbb{R}^{+}$, if
$|U(\mathcal{S}\cup\{z\})-U(\mathcal{S})|\leq\beta(s),$
for all $s\in\mathbb{N}$, $\mathcal{S}\in\mathcal{Z}^{s-1}$ and $z\in\mathcal{Z}$.
Theorem 3.
Suppose Assumptions 1 and 2 hold, and $U$ is $\beta(s)$-deletion stable with $\beta(s)=O(s^{-1})$. Let $\alpha_{0}=s_{0}/n$. For any $\epsilon,\delta\in(0,1)$, set $\bar{s}\asymp\epsilon^{-1/\upsilon}$, $m_{1}\asymp\epsilon^{-(4+2\upsilon)/\upsilon}\log(n/\delta)$, and $m_{2}\asymp\max\{1,\epsilon^{-2}(\alpha_{0}(1-\alpha_{0}))^{2}\log^{3}(n/\delta)\}$. Algorithm 1 gives an $(\epsilon,\delta)$-approximation of $\operatorname{FGSV}(S_{0})$ with $O\left(n\left[1+(\alpha_{0}(1-\alpha_{0}))^{2}(\log n)^{3}\right]\right)$ utility evaluations.
The assumption of $\beta(s)=O(1/s)$ deletion stability in Theorem 3 is mild. Comparable assumptions have appeared in the literature of individual Shapley [19, 36, 37] and other learning problems [1, 9].
Numerous recent advances have developed efficient algorithms for approximating individual Shapley values [13, 34, 24, 26, 3, 27, 23, 38]. These methods target an $(\epsilon,\delta)$-approximation on the full individual Shapley-value vector $\|\widehat{\theta}-\theta\|_{2}$, where $\theta:=(\operatorname{SV}(1),\ldots,\operatorname{SV}(n))$ and $\widehat{\theta}$ is the approximation for $\theta$. The SOTA method achieves this guarantee in $O(n\epsilon^{-2}\log(n/\delta))$ utility evaluations. Approximating $\operatorname{FGSV}(S_{0})=\sum_{i\in S_{0}}\theta_{i}$ by simply summing up individual approximations can lead to an additive error bounded by $\sqrt{s_{0}}\|\widehat{\theta}-\theta\|_{2}$. Thus, achieving the same $(\epsilon,\delta)$-approximation on $\operatorname{FGSV}(S_{0})$ requires tightening the error tolerance to $\epsilon/\sqrt{s_{0}}$, increasing the sample complexity to $O\big{(}\alpha_{0}n^{2}\mathrm{Poly}(\log n)\big{)}$. By directly targeting the group objective, our method avoids this quadratic blow-up and offers a significant speed-up over the SOTA.
3.4 Utility function values for small input sets
In some machine learning scenarios, $U(S)$ might not be well-defined for small $|S|$. For instance, methods such as LLM’s would only produce meaningful result from sufficiently large data sets. Meanwhile, Shapley value emphasizes the contributions from small games, therefore, we cannot simply ignore the small $|S|$ terms in the computation of FGSV (7). One way is to use variants of Shapley value, such as beta-Shapley and Banzhaf, that down-weight small $|S|$ terms, but these alternatives do not satisfy all axioms for faithful group data valuation per Theorem 1. The other way is to fill in $U(S)$ for small $S$ with random or zero values. However, these ad-hoc solutions lack principle and may risk significantly distorting the valuation.
To motivate our approach, we make two simple observations. First, when $S=\emptyset$, the trained method should behave as if it were trained on pure noise, corresponding to a baseline utility value. Second, big-data reliant methods might not value small $S$ very differently than baseline, regardless of their content. In this context, both quality and quantity of data matter for valuation.
Our remedy is very simple. We set a threshold for input size, denoted by $B$, such that $U(S)$ is always well-defined for $|S|\geq B$. If $|S|<B$, we inject $B-|S|$ non-informative data points, to elevate the input size to $B$; otherwise no change is made to $U(S)$. Clearly, we can expect that as the size of informative data (i.e., $|S|$) increases, less amount of non-informative data will be injected, and $U(S)$ gradually becomes non-baseline. For more implementation details, see Algorithm 2 in Appendix B.
While our approach was inspired by the feature deletion technique in machine unlearning, to our best knowledge, we are the first to adapt the idea for data Shapley, in a distinct fashion.
3.5 Examples of second-order stable algorithms
Among the conditions our theory needs, Assumption 2 is arguably the least intuitive. To demonstrate that Assumption 2 is in fact quite mild, here, we verify that it holds for two important applications.
3.5.1 Stochastic Gradient Descent (SGD)
SGD on a training set $\mathcal{S}=\{z_{i}\}_{i=1}^{s}$ minimizes the empirical loss ${\cal L}(w)=\frac{1}{s}\sum_{i=1}^{s}\ell(w;z_{i}).$ Starting from an initial $w_{0}$, SGD runs for $T$ steps, each step $t\in[T]$ updating $w_{t}=w_{t-1}-\alpha_{t}\cdot\frac{1}{m}\sum_{i\in I_{t}}\nabla\ell(w_{t-1};z_{i})$ with learning rate $\alpha_{t}$, where the mini-batch $I_{t}$ of size $m$ is drawn independently and uniformly from $[n]$. Denote the final output by $w_{T}=w(z_{I_{1}},\dots,z_{I_{T}})$. The performance of $w_{T}$ can be represented as some scalar function $u(\cdot)$ (e.g., classification accuracy on a test data). Define the utility as the expected performance:
$U(\mathcal{S}):=$ $~{}\mathbb{E}_{I_{1},\ldots,I_{T}}[u(w_{T})].$ (12)
Assumption 3.
Suppose $\ell$ and $u$ satisfy the following regularity conditions. All conditions hold for some constants $C,L,\beta,\rho>0$ and all $w,w^{\prime},z$.
  1. 1.
    (Smoothness of $u$) $\|\nabla u(w)\|\leq C$ and $\|\nabla u(w^{\prime})-\nabla u(w)\|\leq C\|w^{\prime}-w\|.$
  2. 2.
    ($L$-Lipschitz of $\ell$) $|\ell(w^{\prime};z)-\ell(w;z)|\leq L\|w^{\prime}-w\|.$
  3. 3.
    ($\beta$-smoothness of $\ell$) $\|\nabla_{w}\ell(w^{\prime};z)-\nabla_{w}\ell(w;z)\|\leq\beta\|w^{\prime}-w\|.$
  4. 4.
    (Smoothness of $\ell$’s Hessian) $\|\nabla^{2}_{w}\ell(w^{\prime};z)-\nabla^{2}_{w}\ell(w;z)\|\leq\rho\|w^{\prime}-w\|.$
Proposition 2.
Set $\alpha_{t}\asymp s^{-\tau_{1}}/t$, $m\asymp s^{\tau_{2}}$, and $T\asymp s^{\tau_{3}}$, where constants $\tau_{1}>0$ and $\tau_{2},\tau_{3}\geq 0$. Suppose $\upsilon:=2\tau_{1}+\tau_{2}-1/2\geq 0$. Under Assumption 3, the utility function (12) satisfies:
[…]
3.5.2 Influence Function (IF)
The influence function (IF) method [8, 18] considers the following regularized empirical loss:
$\widehat{\theta}_{\mathcal{S}}:=$ $~{}\arg\min_{\theta}\mathcal{L}(\theta;\mathcal{S}):=\arg\min_{\theta}\Big{\{}s^{-1}\sum_{z\in\mathcal{S}}\ell(\theta;z)+(\lambda/2)\cdot\|\theta\|_{2}^{2}\Big{\}}.$ (13)
Like in Section 3.5.1, suppose the utility can be written as $U(S)=u(\widehat{\theta}_{S})$. Under mild conditions (see Proposition 3), standard IF theory [8, 18] implies that
$U(\mathcal{S}\cup\{z_{1},z_{1}^{\prime}\})-U(\mathcal{S})\approx$ $~{}s^{-1}\nabla_{\theta}u(\widehat{\theta}_{\mathcal{S}})^{\top}H_{\widehat{\theta}_{\mathcal{S}}}^{-1}\Big{\{}\nabla\ell(\widehat{\theta}_{\mathcal{S}};z_{1})+\nabla\ell(\widehat{\theta}_{\mathcal{S}};z_{1}^{\prime})\Big{\}}$ (14)
for any $z_{1},z_{1}^{\prime}$, where $H_{\widehat{\theta}_{\mathcal{S}}}$ is the Hessian of ${\cal L}$ evaluated at $\widehat{\theta}_{\mathcal{S}}$. This yields the following proposition.
Proposition 3.
Suppose $u$ is continuously differentiable with $L$-Lipschitz and bounded gradient. Assume that the loss function $\ell(\theta;z)$ is convex in $\theta$, three times continuously differentiable with all derivatives up to order three uniformly bounded. Then for any $z_{1},z_{1}^{\prime},z_{2},z_{2}^{\prime}\in\mathcal{Z}$,
$\big{|}U(\mathcal{S}\cup\{z_{1},z_{1}^{\prime}\})-U(\mathcal{S}\cup\{z_{1},z_{2}\})-U(\mathcal{S}\cup\{z_{1}^{\prime},z_{2}^{\prime}\})+U(\mathcal{S}\cup\{z_{2},z_{2}^{\prime}\})\big{|}\lesssim s^{-2}.$ (15)
Proposition 3 is not surprising in view of (14), as the main terms cancel out on the LHS of (15).
4 Experiments
We empirically compare our method (FGSV) to the approach of first computing individual Shapley values using SOTA methods and then summing them to form a group valuation. The results demonstrate the superiority of our method in both speed and approximation accuracy and its faithfulness under shell company attack in applications such as copyright attribution in generative AI and explainable AI. We report the key findings in the main paper and relegate full details to Appendix A.
4.1 Approximation accuracy and computational efficiency
In this experiment, we compare our method to the following benchmarks: (1) Permutation-based estimator [2, 6] that averages marginal contributions over random data permutations; (2) Group Testing-based estimator [12, 34] that uses randomized group inclusion tests to estimate pairwise Shapley value differences; (3) Complementary Contribution estimator [38] that uses stratified sampling of complementary coalitions; (4) One-for-All estimator [24] that uses weighted subsets and reuses utility evaluations; (5) KernelSHAP [26], a regression-based method with locally weighted samples; (6) Unbiased KernelSHAP [3], KernelSHAP with a bias-correction; and (7) LeverageSHAP [27] that speeds up KernelSHAP by weighted sampling based on leverage scores.
Experimental setup.
We consider the Sum-of-Unanimity (SOU) cooperative game [24], where the individual Shapley values have a closed-form expression. Specifically, the utility function is defined as $U(S)=\sum_{j=1}^{d}\alpha_{j}\mathbf{1}_{\mathcal{A}_{j}\subseteq S}$, and the corresponding Shapley value for player $i\in[n]$ is given by $\operatorname{SV}(i)=\sum_{j=1}^{d}\frac{\alpha_{j}}{|\mathcal{A}_{j}|}\mathbf{1}_{i\in\mathcal{A}_{j}}$. Here, each subset $\mathcal{A}_{j}$ is generated by sampling a random size uniformly from $1$ to $n$ and then drawing that many players without replacement. We set $d=n^{2}$ and the coefficient $\alpha_{j}$ as the average of the weights of all players $i\in\mathcal{A}_{j}$, where each player $i$ has a weight $\frac{(i\bmod 4)}{4}$. We partition all players into $4$ groups based on their indices: $S_{k}=\big{\{}i\in[n]:i\equiv k-1(\bmod 4)\big{\}}$ for $k\in[4]$, and our goal is to estimate the group value $\operatorname{FGSV}(S_{k})=\sum_{i\in S_{k}}\operatorname{SV}(i)$. Given a fixed budget of 20,000 utility function evaluations, we record the absolute relative error of the $\operatorname{FGSV}$ estimate every 200 iterations for each method. We summarize the performance using the Area Under the Convergence Curve (AUCC): $\operatorname{AUCC}(S_{k})=\frac{1}{100}\sum_{t=1}^{100}|\frac{\operatorname{FGSV}(S_{k})-\widehat{\operatorname{FGSV}}^{(200\cdot t)}(S_{k})}{\operatorname{FGSV}(S_{k})}|,$ where $\widehat{\operatorname{FGSV}}^{(\tau)}(\cdot)$ is the estimate after $\tau$ iterations. Lower AUCC indicates better accuracy and faster convergence. We also record the average runtime per iteration for each method.
Figure 2: Performance comparison in the SOU game. Top: Our method (FGSV) achieves the lowest AUCC across all problem sizes. Bottom: Our method costs the lowest runtime per iteration.
Results.
Figure 2 summarizes the average performance over $30$ replications for $n\in\{64,128,256\}$. The top row shows that our method consistently achieves the lowest AUCC across all problem sizes $n$. The bottom row suggests that, while using the same utility evaluation budget, our method is among the fastest. In contrast, baselines such as Group Testing, KernelSHAP, and Unbiased KernelSHAP incur substantial higher computational overhead, due to internal optimization processes. Overall, our method achieves both faster convergence and improved computational efficiency, which demonstrates the benefits of directly estimating the group value rather than aggregating individual Shapley estimates.
4.2 Application to faithful copyright attribution in generative AI
Group data valuation is important for fairly compensating copyright holders whose data are used to train generative AI models. Existing approaches, such as the Shapley Royalty Share (SRS) proposed by [35], adopt the GaI approach based on GSV. Given a partition of the training data into $K$ disjoint groups $S_{1},\dots,S_{K}$, SRS is defined as $\operatorname{SRS}(S_{k};x_{\mathrm{gen}}):=\frac{\operatorname{GSV}(S_{k};x_{\mathrm{gen}})}{\sum_{j=1}^{K}\operatorname{GSV}(S_{j};x_{\mathrm{gen}})}$. However, as discussed in Section 2, GSV is prone to the shell company attack. To address this vulnerability, we propose the Faithful Shapley Royalty Share (FSRS) that replaces GSV with our $\operatorname{FGSV}$ to faithfully reward individuals in a group by their contributions, not by tactical grouping strategies:
$\operatorname{FSRS}(S_{k};x_{\mathrm{gen}}):=\frac{\operatorname{FGSV}(S_{k};x_{\mathrm{gen}})}{\sum_{j=1}^{K}\operatorname{FGSV}(S_{j};x_{\mathrm{gen}})}.$
Experimental setup.
Following [35], we fine-tune Stable Diffusion v1.4 [28] using Low-Rank Adaptation (LoRA; [11]) on four brand logos from FlickrLogo-27 [15]. The utility $U(\cdot;x^{(\mathrm{gen})})$ is the average log-likelihood of generating $20$ brand-specific images $x^{(\mathrm{gen})}$ using the prompt “A logo by [brand name]” (see example images in Panel (a) of Figure 3). We compare SRS and FSRS under two grouping scenarios: (1) $30$ images from each brand form a single group, and (2) the Google and Sprite datasets are each split into two subgroups (20/10 images), launching a shell company attack. Importantly, the total data per brand remains unchanged across two scenarios.
Figure 3: Comparison of SRS and FSRS for copyright attribution. (a) Example images generated using brand prompts. (b) Shapley Royalty Share (SRS, [35]) based on GSV. (c) Faithful SRS (FSRS, our method) based on $\operatorname{FGSV}$. Blue bars: valuation under Scenario 1 (single group per brand); orange bars: valuation under Scenario 2 (Google/Sprite data each split into size-20/10 subgroups, colored in dark and light orange).
Results.
Figure 3 illustrates the impact of the shell company attack on copyright attribution under SRS and FSRS. Each subplot corresponds to a brand-specific prompt and shows the royalty shares assigned to each brand. Conceptually, the brand corresponding to the prompt should receive the highest royalty share. Panel (b) shows that applying a shell company attack that favors the Google and Sprite groups indeed inflates their total SRS shares substantially, while reducing others’, despite the contents contributed by each group remain unchanged. This produces misleading results—for example, Google and Sprite receive higher overall SRS than Vodafone under the “A logo by Vodafone” prompt. In contrast, Panel (c) shows that FSRS yields consistent and stable valuations under the shell company attack. This demonstrates that FSRS mitigates the effects of strategic data partitioning and provides a more faithful reflection of group contribution for copyright attribution.
4.3 Application to faithful explainable AI
Group data valuation is also a crucial tool in explainable AI, providing interpretable summaries of data contributions at the group level— particularly in contexts where groups correspond to socially or scientifically significant categories. For example, [6] used GSV to quantify the contributions of data from different demographic groups to patient readmission prediction accuracy [31]. However, GSV’s sensitivity to the choice of grouping can cause group values to fluctuate dramatically under different data partitions, leading to inconsistent interpretations. To address this problem, we recommend practitioners to use FGSV and empirically demonstrate that it produces more consistent and reliable interpretations across a variety of grouping configurations.
Experimental setup.
We conduct our experiment on the Diabetes dataset [5], which contains 442 individuals, each described by 10 demographic and health-related features (e.g., sex, age, and BMI). The task is to predict the progression of diabetes one year after baseline. We construct 7 grouping schemes, partitioned by all non-empty subsets of the variables sex, age, and BMI, excluding the trivial no-partition case. For the continuous variables age and BMI, we discretize each into three quantile-based categories. Overall, we will have between 2 to 18 total groups. For each grouping scheme, we compute GSV exactly and estimate FGSV via 30 Monte Carlo replications. Our predictive model is ridge regression, and we measure utility as the negative mean squared error on a held-out test set, with the null utility set to the variance of the test responses. Thus, higher group valuations indicate greater contributions to the model’s predictive accuracy. To compare category-level contributions across different grouping schemes, we aggregate group values as follows: for a given variable (e.g., sex), we sum the (F)GSVs of all groups that include that category. This allows for consistent cross-scheme comparisons, where one would expect a category to receive stable valuations after aggregation across different grouping strategies if the valuation method is robust.
Figure 4: Comparison of GSV (top row) and FGSV (bottom row) in a regression task for explainable AI. Each column aggregates category-level values for a specific variable: sex (left), age (middle), and BMI (right). Shaded areas represent ±$\pm 1$ standard deviation across 30 replications.
Results.
Figure 4 visualizes the aggregated category-level values from GSV (top row) and FGSV (bottom row) across various grouping schemes. FGSV produces significantly more stable and consistent group rankings compared to GSV. For instance, in the age-wise plots (second column), GSV gives inconsistent rankings: each of the three age groups—“Young,” “Middle,” and “Old”—appears as the top-ranked group under at least one grouping scheme. This instability undermines the reliability of the interpretation. In contrast, FGSV consistently assigns the highest value to the “Young” group across all grouping schemes, demonstrating its robustness. Overall, FGSV provides more reliable explanations that are less sensitive to arbitrary grouping schemes.
5 Conclusion and discussion
In this paper, we proposed FGSV for faithful group data valuation. We showed that FGSV is the unique group valuation method that satisfies a desirable set of principles, including faithfulness, which ensures that group value remains unchanged to arbitrary re-grouping among other players, thereby defending against the shell company attack. Our algorithm also achieves lower sample complexity over SOTA methods that sum up individual Shapley value estimates, as demonstrated through both theoretical analysis and numerical experiments. We further illustrated the robustness of FGSV in applications to copyright attribution and explainable AI, where it faithfully reflects and fairly rewards individual contributions from group members.
Beyond the shell company attack, there exists another unfair competition strategy, namely, the copier attack, in which a group may steal valuation from other groups by duplicating their high-value data points. While a principled defense against this attack remains an open problem, a straightforward ad-hoc remedy is to pre-process the dataset to detect and remove such duplicated entries before applying our FGSV method.
Code
The code and instructions to reproduce the experiments are provided in the supplementary material and available at https://github.com/KiljaeL/Faithful_GSV.
Acknowledgements
Lee and Zhang were supported by NSF DMS-2311109. Liu and Tang were supported by NSF DMS-2412853.
Appendix A Experimental Details
This appendix outlines the implementation details for all experiments in the main paper. To evaluate the effectiveness, efficiency, and robustness of FGSV, four experimental settings were employed. These consist of: (i) a synthetic data example in the introduction to illustrate the shell company attack (Section 1); (ii) a benchmark comparison on the SOU cooperative game (Section 4.1); (iii) an application to copyright attribution using a generative AI model with FlickrLogo-27 Dataset (Section 4.2); and (iv) an application to explainable AI using the Diabetes dataset (Section 4.3).
All experiments were conducted on the Unity high-performance computing cluster, provided by the College of Arts and Sciences at The Ohio State University. The copyright attribution experiment (Section 4.2) was executed on a GPU node equipped with an NVIDIA V100 GPU, 32GB VRAM, whereas all other experiments were conducted on CPU nodes featuring Intel Xeon E5-2699 v4 processors and 256 GB RAM.
A.1 Motivational Example (Section 1)
A synthetic setting is constructed to illustrate the shell company attack and its effect on group-level data valuation. We generate synthetic data for binary classification from a Gaussian mixture. Specifically, given a class label $y_{i}\in\{0,1\}$, each sample $x_{i}\in\mathbb{R}^{2}$ is drawn from $\mathcal{N}(\mu_{1},I_{2})$ if $y_{i}=0$, and from $\mathcal{N}(\mu_{2},I_{2})$ if $y_{i}=1$, where $\mu_{1}=[-3,0]^{\top}$ and $\mu_{2}=[3,0]^{\top}$. The class label is sampled independently with equal probability $p=0.5$. An additional $n=200$ samples are independently generated from the same distribution for testing.
The training data is partitioned into two equal-sized groups of 100 samples each. To simulate the shell company attack, the second group is further subdivided evenly into 2, 3, or 4 subgroups. This results in a total of $k\in\{2,3,4,5\}$ disjoint groups: the first group consistently contains 100 samples, while the remaining $k-1$ groups are allocated approximately $100/(k-1)$ samples each, with any remainder distributed to maintain balance (e.g., 33, 33, and 34 when $k=4$).
For each subset $S$ of the training data, a logistic regression classifier is trained. The utility function $U$ is defined as its classification accuracy on a held-out test set. If $S$ contains only a single class, the utility is set to a constant value of $1/2$, corresponding to the expected accuracy of a random classifier.
While the experiment is qualitative and does not include confidence intervals, repeated runs with different random seeds produced consistent trends. FGSV is estimated using Algorithm 1 with threshold parameter $\bar{s}=20$ and Monte Carlo sample sizes $m_{1}=m_{2}=2000$.
A.2 Comparison with benchmark methods (Section 4.1)
For all methods compared here, we adjust their configurations, such that the total number of utility evaluations is aligned to the fixed number of 20,000. Note that since FGSV estimates the value for one group at a time, for fairness, the total budget is equally divided among the four groups, with 5,000 utility evaluations allocated per group. For FGSV, we set the threshold parameter to $\bar{s}=10$ and choose $m_{1}=m_{2}$ such that the total number of utility function calls sums to 20,000. Below, we describe the setup and computational structure of each method in detail, largely based on the implementation and descriptions from [24].
  • Permutation estimator [2, 6].
    The permutation estimator computes Shapley values by averaging marginal contribution across random permutations. At each iteration, it draws a permutation $\pi$ uniformly at random from all possible orderings of the $n$ elements and computes a Monte Carlo estimate based on the following alternative representation of the individual Shapley value:
    $\operatorname{SV}(i)=\mathbb{E}_{\pi}\left[U(S_{\pi,i}\cup\{i\})-U(S_{\pi,i})\right],$
    where $S_{\pi,i}$ denotes the set of indices that appear before $i$ in permutation $\pi$. The final estimate is obtained by averaging the marginal contributions over $T$ sampled permutations:
    $\widehat{\operatorname{SV}}_{\text{perm}}(i)=\frac{1}{T}\sum_{t=1}^{T}\left[U(S_{\pi_{t},i}\cup{i})-U(S_{\pi_{t},i})\right],$
    where $\pi_{t}$ denotes the $t$-th sample of random permutation. To simultaneously estimate $\operatorname{SV}(i)$ for all data points $i\in[n]$ per permutation, this method requires $n+1$ utility evaluations – one for the empty set $U(\varnothing)$, and one each within incremental updates, $U(\{\pi(1),\dots,\pi(j)\})$ for $j=1,\dots,n$. Therefore, we set $T=\lceil\frac{20000}{n+1}\rceil$ to maintain the total utility evaluation budget. For the last sampled permutation, the estimator stopped when the number of evaluations reached exactly 20,000.
  • Group Testing estimator [12, 34].
    This estimator computes pairwise differences of Shapley values rather than individual values. At each step, it draws a random size of subset $s\in[n]$ with $p(s)\propto\frac{1}{s}+\frac{1}{n-s+1}$, and draws a random subset $S_{t}\subseteq[n+1]$ of size $s$, where $(n+1)$-th element represents a dummy player, i.e., a player whose contribution is always zero and serves as a reference for baseline utility. In particular, for a given number of utility evaluations, $T$, a matrix $B\in\mathbb{R}^{T\times(n+1)}$ is constructed, where the $(t,i)$-th entry is given by:
    $B_{t,i}=\begin{cases}U(S_{t}\setminus\{n+1\})&\text{if }i\in S_{t},\\ 0&\text{otherwise}.\end{cases}$
    We set $T=20000$. Then, the Shapley value is approximated by
    $\widehat{\operatorname{SV}}_{\text{gt}}(i)=\frac{Z}{T}\sum_{t=1}^{T}\left(B_{t,i}-B_{t,n+1}\right),$
    where $B_{t,i}$ denotes the utility observed at iteration $t$ when player $i\in S_{t}$, and $Z=\sum_{s=1}^{n}p(s)$ is a normalization constant.
  • Complement Contribution estimator [38].
    The complement estimator approximates Shapley values by leveraging the symmetry between a subset and its complement. In particular, it relies on another equivalent representation of the Shapley value:
    $\operatorname{SV}(i)=\frac{1}{n}\sum_{S\subseteq[n]\setminus\{i\}}\frac{U(S\cup\{i\})-U([n]\setminus(S\cup\{i\}))}{\binom{n-1}{|S|}}.$
    This expression enables the estimator to reuse utility evaluations for both $S$ and its complement $[n]\setminus S$. At each iteration, it samples a subset size $s\in[n]$ uniformly and then uniformly samples a subset $S\subseteq[n]$ of size $s$. Then, the final estimator is:
    $\widehat{\operatorname{SV}}_{\text{cc}}(i)=\frac{1}{n}\sum_{s=1}^{n}\frac{1}{T_{i,s}}\sum_{t=1}^{T}\left[v_{t}\left\{\mathbb{I}(i\in S_{t},|S_{t}|=s)-\mathbb{I}(i\notin S_{t},|S_{t}|=n-s)\right\}\right],$
    where $v_{t}=U(S_{t})-U([n]\setminus S_{t})$ and $T_{i,s}$ is the number of such samples satisfying one of the two conditions. In our experiment, we set $T=20000/2=10000$ so that each iteration, which requires two utility evaluations, results in exactly 20,000 total evaluations.
  • One-for-All estimator [24].
    The One-for-All estimator is based on the following alternative formulation of the Shapley value [^1] While the One-for-All estimator can estimate a variety of values (e.g., Banzhaf [33], Beta-Shapley [19]), we focus on its use for Shapley value estimation.:
    $\operatorname{SV}(i)=\frac{1}{n}\sum_{s=1}^{n}\left(\mathbb{E}_{S\subseteq[n],|S|=s,i\in S}[U(S)]-\mathbb{E}_{S\subseteq[n],|S|=s-1,i\notin S}[U(S)]\right),$
    This formulation allows each sampled subset to contribute to the estimates of all players, enabling efficient sample reuse. First, it deterministically allocates $22n+$ 2 utility evaluations for subset sizes $s\in\{0,1,n-1,n\}$, and uses them to compute the corresponding expectations exactly. Then, for the remaining possible subset sizes $\{2,\dots,n-2\}$, it samples $s$ with a predefined sampling probability $q(s)$, and then samples a subset $S_{t}\subseteq[n]$ of size $s$ uniformly at random. The final estimation is:
    $\widehat{\operatorname{SV}}_{\text{ofa}}(i)$ $=\frac{1}{n}(U([n])-U(\varnothing))$
    $+\frac{1}{n(n-1)}\left(\sum_{i\in S,|S|=n-1}U(S)-\sum_{i\notin S,|S|=1}U(S)\right)$
    $+\frac{1}{n}\sum_{s=2}^{n-2}\left(\frac{1}{T_{i,s}^{\text{in}}}\sum_{t=1}^{T}U(S_{t})\cdot\mathbb{I}(|S_{t}|=s,i\in S_{t})-\frac{1}{T_{i,s}^{\text{out}}}\sum_{t=1}^{T}U(S_{t})\cdot\mathbb{I}(|S_{t}|=s,i\notin S_{t})\right),$
    where $T_{i,s}^{\text{in}}$ and $T_{i,s}^{\text{out}}$ denote the number of Monte Carlo samples of size $s$ in which $i$ is included and excluded, respectively. In our experiment, we used $q(s)\propto\frac{1}{\sqrt{s(n-s)}}$, which is proved to be the optimal sampling distribution for Shapley value estimation. Next, after allocating $22n+$ 2 evaluations deterministically for subset sizes $s\in\{0,1,n-1,n\}$, the remaining evaluations are used for Monte Carlo sampling. That is, we set $T=20000-(2n+2)$ for sampling subset sizes $s\in\{2,\dots,n-2\}$.
  • KernelSHAP [26].
    The Shapley value can be characterized as the solution to the following constrained optimization problem:
    $\operatorname{SV}=\arg\min_{\phi\in\mathbb{R}^{n}}\sum_{\varnothing\subsetneq S\subsetneq[n]}w_{S}\left(U(S)-U(\varnothing)-\sum_{i\in S}\phi_{i}\right)^{2},\quad\text{subject to}\quad\sum_{i=1}^{n}\phi_{i}=U([n])-U(\varnothing),$
    where $w_{S}\propto\frac{1}{\binom{n-2}{|S|-1}}$ is the kernel weight assigned to each subset $S$. Based on this formulation, KernelSHAP constructs a sample-based approximation of the objective and solves the resulting weighted least squares problem to obtain an estimate of the Shapley value. In particular, at each iteration, it samples a subset size $s\in\{1,\dots,n-1\}$ according to the distribution $p(s)\propto\frac{1}{s(n-s)}$, and then draws a subset $S_{t}\subseteq[n]$ of size $s$ uniformly at random. Each subset is encoded as a binary indicator vector $\mathbf{1}_{S}\in\{0,1\}^{n}$, and the following quantities are updated:
    $\widehat{A}=\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}_{S_{t}}\mathbf{1}_{S_{t}}^{\top},\quad\widehat{b}=\frac{1}{T}\sum_{t=1}^{T}\left(U(S_{t})-U(\varnothing)\right)\cdot\mathbf{1}_{S_{t}}.$
    The KernelSHAP estimator is obtained in closed form as:
    $\widehat{\operatorname{SV}}_{\text{ks}}=\widehat{A}^{-1}\left(\widehat{b}-\frac{\mathbf{1}^{\top}\widehat{A}^{-1}\widehat{b}-U([n])+U(\varnothing)}{\mathbf{1}^{\top}\widehat{A}^{-1}\mathbf{1}}\cdot\mathbf{1}\right),$
    where $\mathbf{1}\in\mathbb{R}^{n}$ denotes the vector with all-one entries. We set $T=20000$ in our implementation of KernelSHAP.
  • Unbiased KernelSHAP [3].
    The Unbiased KernelSHAP estimator is a variant of KernelSHAP that utilizes the fact that the exact gram matrix, $A=\mathbb{E}[\mathbf{1}_{S}\mathbf{1}_{S}^{\top}]$, admits a closed-form expression under the same distribution $w_{S}\propto\frac{1}{\binom{n-2}{|S|-1}}$. Instead of estimating $A$ empirically from samples, this estimator directly computes $A$, where its $(i,j)$-entry is defined as:
    $A_{ij}=\begin{cases}\frac{1}{2}&\text{ if }i=j,\\ \frac{1}{n(n-1)}\frac{\sum_{s=2}^{n-1}\frac{s-1}{n-s}}{\sum_{s=1}^{n-1}\frac{1}{s(n-s)}}&\text{otherwise.}\end{cases}$
    Unbiased KernelSHAP replaces the empirical matrix $\widehat{A}$ in KernelSHAP with its closed-form expectation $A$:
    $\widehat{\operatorname{SV}}_{\text{uks}}=A^{-1}\left(\widehat{b}-\frac{\mathbf{1}^{\top}A^{-1}\widehat{b}-U([n])+U(\varnothing)}{\mathbf{1}^{\top}A^{-1}\mathbf{1}}\cdot\mathbf{1}\right).$
    In our implementation, we set $T=20000$.
  • LeverageSHAP [27].
    LeverageSHAP is another variant of KernelSHAP designed to reduce the variance of the estimator and improve computational efficiency. Unlike KernelSHAP, which samples subset sizes $s\in\{1,\dots,n-1\}$ with the weighting $p(s)\propto\frac{1}{s(n-s)}$, LeverageSHAP draws $s$ uniformly (i.e., $p(s)\propto 1$) and compensates for the mismatch via a correction factor $w(s)=\sqrt{s(n-s)}$. Also, it adopts paired sampling, which also includes complement $\bar{S}_{t}=S_{t}^{c}$ in the sample pool when each subset $S_{t}$ is sampled. Based on these modification, it computes the following quantities:
    $\widetilde{A}$ $=\frac{1}{2T}\sum_{t=1}^{T}w(|S_{t}|)\left\{\mathbf{1}_{S_{t}}\mathbf{1}_{S_{t}}^{\top}+\mathbf{1}_{\bar{S}_{t}}\mathbf{1}_{\bar{S}_{t}}^{\top}\right\}$
    $\widetilde{b}$ $=\frac{1}{2T}\sum_{t=1}^{T}w(|S_{t}|)\left\{\left(U(S_{t})-U(\varnothing)\right)\cdot\mathbf{1}_{S_{t}}+\left(U(\bar{S}_{t})-U(\varnothing)\right)\cdot\mathbf{1}_{\bar{S}_{t}}\right\}.$
    The final LeverageSHAP estimator is then computed by:
    $\widehat{\operatorname{SV}}_{\text{ls}}=\widetilde{A}^{-1}\left(\widetilde{b}-\frac{\mathbf{1}^{\top}\widetilde{A}^{-1}\widetilde{b}-U([n])+U(\varnothing)}{\mathbf{1}^{\top}\widetilde{A}^{-1}\mathbf{1}}\cdot\mathbf{1}\right).$
    For LeverageSHAP, we set $T=20000/2=10000$ since each iteration requires two utility evaluations due to the use of paired sampling.
Although all methods use the same total number of utility evaluations (20,000), their actual runtimes differ substantially, as shown in the second row of Figure 2. These discrepancies are attributable to several algorithmic and implementation-specific factors, including computational complexity and sampling distribution over subset sizes.
First, Group Testing requires the construction of a utility matrix $B\in\mathbb{R}^{T\times(n+1)}$. Maintaining this large matrix introduces relatively high memory usage and per-iteration computational overhead, as each utility evaluation must be copied into multiple locations with structural indexing. In contrast, other estimators maintain only low-dimensional accumulators or regression statistics, which impose negligible memory cost.
Second, the cost of evaluating $U(S)$ varies dramatically with the size of the subset, $|S|=s$. Consequently, the subset size sampling distribution required by each method significantly impacts speed. Among the benchmarks, Group Testing: $p(s)\propto\frac{1}{s}+\frac{1}{n-s+1}$; One-for-All: $q(s)\propto\frac{1}{\sqrt{s(n-s)}}$; and both KernelSHAP and Unbiased KernelSHAP: $p(s)\propto\frac{1}{s(n-s)}$. They all employ non-uniform subset size distributions that tend to oversample either very small or very large subsets. This matters significantly in the SOU game, where evaluating $U(S)$ requires checking whether $\mathcal{A}_{j}\subseteq S$ holds for all $j\in[d]$. Figure 5 empirically demonstrates the near-exponential growth in computation time for evaluating $U(S)$ as a function of subset size $s$, across varying values of $n\in\{64,128,256\}$. This occurs because the time required for set inclusion checks (e.g., issubset() in Python) increases rapidly with subset size. As a result, estimators that disproportionately sample such extreme subset sizes tend to exhibit higher average computation time per evaluation compared to those that sample subset sizes uniformly (e.g., Permutation, Complement Contribution, LeverageSHAP). Thus, sampling behavior—not just the number of evaluations—directly impacts computational efficiency. On the other hand, FGSV does not sample subset sizes $s$, but instead performs estimation through an explicit loop over all possible values of $s$. As shown in Algorithm 2, when $s<\bar{s}$, it evaluates the utility function over all grid points $s_{1}\in[\max\{0,s+s_{0}-n\},\min\{s,s_{0}\}]$. For $s\geq\bar{s}$, the approximation is performed using only two representative values of $s_{1}$, thereby significantly reducing the number of required evaluations. This thresholding mechanism ensures that, under a fixed budget of utility calls, FGSV concentrates more computation on smaller subset sizes, leading to lower overall runtime.
Figure 5: Empirical average runtime (in seconds) of $U(S)$ evaluation as a function of subset size $s=|S|$. Each curve represents the mean over 50 randomly sampled subsets of size $s$; shaded areas indicate ±$\pm 1$ standard deviation.
Third, KernelSHAP and its variants—Unbiased KernelSHAP and LeverageSHAP—require solving a constrained least squares problem involving dense matrix inversions and multiplications. These linear algebra operations, performed either during or after sampling, dominate the overall computational cost, making these methods significantly slower than alternatives that rely solely on running averages or marginal contribution estimates.
A.3 Faithful copyright attribution in generative AI (Section 4.2)
We evaluate SRS and FSRS in the context of group-level copyright attribution for generative models, using a logo generation task based on Stable Diffusion. To facilitate comparison, we mostly followed the experimental setup in [35]. Nonetheless, for clarity and completeness, we summarize the relevant details below.
To construct the experiment, we select four logo classes (Google, Sprite, Vodafone, and Starbucks) for evaluation, and use the remaining 23 brands to initialize a baseline model via fine-tuning of Stable Diffusion v1.4 [28]. We then assess the impact of fine-tuning with each of the four excluded brands by measuring how their inclusion alters the generation behavior of the model.
Fine-tuning is performed using Low-Rank Adaptation (LoRA) [11], a technique that inserts trainable low-rank matrices into the attention layers of the diffusion model. The rank $r$ and scaling factor $\alpha$ are both set to 8, and fine-tuning is only applied to the attention layers of the UNet module, while all other weights of the Stable Diffusion model are kept frozen. This configuration ensures efficient and stable adaptation to the small-scale dataset, in line with standard LoRA usage. [^2] https://huggingface.co/blog/lora
We use a learning rate of 0.0001, a batch size of 4, and train for 10 epochs. After fine-tuning, we generate $N_{\rm MC}=20$ images for each of the four selected brands using the prompt “A logo by [brand name].” Image generation is performed with 25 denoising steps and classifier-free guidance (scale 7.5), implemented using the DDPMScheduler.
Following the formulation of [35], we define the utility function as the log-likelihood of generated images. For a given length of the denoising steps $T$, let $(x_{T},\dots,x_{1},x_{0})$ denote the reverse trajectory defined by the diffusion scheduler, where $x_{T}\sim\mathcal{N}(0,I)$ is the initial latent noise, and $x_{0}=x_{(\text{gen})}$ is the final generated image. The likelihood of $x_{(\text{gen})}$ under model parameters $\theta$ is defined as:
$p_{\theta}(x_{\text{(gen)}})$ $=p_{\theta}(x_{0})=\mathbb{E}_{x_{1}}\left[p_{\theta}(x_{0}\mid x_{1})\right],$
by the Markov property of the diffusion process. The conditional distribution $p_{\theta}(x_{0}\mid x_{1})$ is Gaussian, given by
$p_{\theta}(x_{0}\mid x_{1})\sim\mathcal{N}\left(x_{0};\frac{1}{\sqrt{\alpha_{1}}}\left(x_{1}^{(j)}-\sqrt{1-\alpha_{1}}\,\widehat{\epsilon}_{\theta}(x_{1}^{(j)},1)\right),\sigma_{1}^{2}I\right),$
where $\widehat{\epsilon}_{\theta}(x_{t},t)$ denotes the predicted noise at step $t$, and $\alpha_{1},\sigma_{1}^{2}$ are scheduler-specific constants. Based on these formulation, the likelihood of $x_{\text{(gen)}}$ can be approximated as
$\frac{1}{N_{\rm MC}}\sum_{j=1}^{N_{\rm MC}}\mathcal{N}\left(x_{0};\frac{1}{\sqrt{\alpha_{1}}}\left(x_{1}^{(j)}-\sqrt{1-\alpha_{1}}\widehat{\epsilon}_{\theta}(x_{1}^{(j)},1)\right),\sigma_{1}^{2}I\right),$
where $x_{1}^{(j)}$ for $j=1,\dots,N_{\rm MC}$ is sampled by reversing the diffusion process from a standard Gaussian noise vector $x_{T}^{(j)}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,I)$.
In our setting, the model is initialized from a pretrained checkpoint based on the remaining 23 brands, which stabilizes the fine-tuning process even when the input subset is small. This design parallels the data augmentation strategy described in Section 3.4, where non-informative examples are added to ensure that the utility function remains well-defined. Ultimately, the resulting setup can be conceptually viewed as one where the utility function consistently receives a sufficiently large dataset as input. Hence, the computation procedure can be understood as effectively a special case of Algorithm 2, with the difference that our computation injects informative, rather than non-informative, data points. As a result, when theoretically understanding the performance of our method here, we resort to Theorem 2, thinking that $s$ is lower-bounded.
Finally, we used $m=2$ for this approximation. This choice was primarily informed by practical constraints: the computation for a single sample, including both fine-tuning and utility evaluation, takes nearly 30 hours given the limited hardware resources available to us. On the other hand, however, we observed that the FSRS estimates well-illustrates the faithfulness property of FGSV—the preservation of group-level value under further partitioning—even within each single experiment. As shown in Figure 3, which displays the result from a single experiment, the FSRS values remained stable before and after splitting the Google and Sprite brands. This suggests that the empirical evidence is clear even with a small $m$. To understand this phenomenon, we attribute such empirical stability to the fact that the diffusion model here was initialized by a pretrained baseline model. The baseline model was trained by on big data with very significant computing resources, which tends to yield rather stable initializations. Consequently, fine-tuning outcomes across random seeds tend to inherit some stability and would not vary wildly.
A.4 Faithful explainable AI (Section 4.3)
We evaluated FGSV and GSV on the Diabetes dataset [5], which includes 442 samples with 10 demographic and health-related features. The dataset is preprocessed by centering and standardizing the response variable and splitting the dataset into 350 training and 92 test samples using a fixed random seed.
For the base model, we use ridge regression with regularization strength $\alpha=0.01$. The utility $U(S)$ is defined as the negative mean squared error (MSE) on the test set; for empty subsets $S$, the utility is defined as the negative variance of the test labels, which corresponds to a baseline predictor that always outputs the test mean. FGSV is estimated using Algorithm 2 with parameters $\bar{s}=35$ and $m_{1}=m_{2}=1000$, while GSV is computed exactly. Each setting is repeated for 30 times using independent Monte Carlo replications.
Appendix B Approximation algorithm with non-informative data augmentation
Algorithm 2 Approximate FGSV with non-informative data augmentation for small input sizes
0:  Dataset $\mathcal{D}$, group $\mathcal{S}_{0}$, size threshold $B$, subsample size $m$, non-informative distribution $\mathcal{P}_{\operatorname{null}}$.
1:  Initialize $n\leftarrow|\mathcal{D}|$, $s_{0}\leftarrow|S_{0}|$, $\alpha_{0}\leftarrow s_{0}/n$.
2:  Initialize total sum for correction terms $\widehat{\mathcal{T}}_{sum}\leftarrow 0$.
3:  for $s=1$ to $n-1$ do
4:     Set $s_{1}^{*}\leftarrow\lfloor s\alpha_{0}\rfloor$. {Expected intersection size for current $s$}
5:     Initialize sum of utility differences $S_{\Delta U}\leftarrow 0$.
6:     for $j=1$ to $m$ do
7:        Sample a base tuple $(\mathcal{S},z_{1},z_{2})$ i.i.d. from $\mathscr{B}_{s,s_{1}^{*}}=\{(\mathcal{S},z_{1},z_{2}):\mathcal{S}\subseteq\mathcal{D},|\mathcal{S}|=s,|\mathcal{S}\cap\mathcal{S}_{0}|=s_{1},z_{1}\in\mathcal{S}_{0}\setminus\mathcal{S},z_{2}\in\mathcal{S}_{0}^{c}\setminus\mathcal{S}\}$, where $\mathcal{S}_{0}=\{z_{i}:i\in S_{0}\}$.
8:        if $s<B$ then
9:           Sample augmentation set $\mathcal{S}_{null}=\{z^{\prime}_{1},\dots,z^{\prime}_{B-s}\}$ where $z^{\prime}_{l}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{P}_{\operatorname{null}}$.
10:           $U_{1}\leftarrow U(\mathcal{S}\cup\{z_{1}\}\cup\mathcal{S}_{null})$.
11:           $U_{2}\leftarrow U(\mathcal{S}\cup\{z_{2}\}\cup\mathcal{S}_{null})$.
12:        else
13:           $U_{1}\leftarrow U(\mathcal{S}\cup\{z_{1}\})$.
14:           $U_{2}\leftarrow U(\mathcal{S}\cup\{z_{2}\})$.
15:        end if
16:        $S_{\Delta U}\leftarrow S_{\Delta U}+(U_{1}-U_{2})$.
17:     end for
18:     $\widehat{\Delta\mu}\left(\frac{s_{1}^{*}}{s};s,s_{0},n\right)\leftarrow\frac{1}{m}S_{\Delta U}$.
19:     $\widehat{\mathcal{T}}(s)\leftarrow\frac{n}{n-1}\alpha_{0}(1-\alpha_{0})\cdot\widehat{\Delta\mu}\left(\frac{s_{1}^{*}}{s};s,s_{0},n\right)$.
20:     $\widehat{\mathcal{T}}_{sum}\leftarrow\widehat{\mathcal{T}}_{sum}+\widehat{\mathcal{T}}(s)$.
21:  end for
22:  $G_{0}\leftarrow\frac{s_{0}}{n}\left[U([n])-U(\varnothing)\right]$.
23:  return  $G_{0}+\widehat{\mathcal{T}}_{sum}$.